Consider a spacecraft moving at 0.95c travelling past an observer on Earth. The said observer and the occupant of the ship each start identical alarm clocks set to ring after 15 hours. With respect to the observer on Earth, at what time will the Earth clock read when the spacecraft clock rings? 


1
Expert's answer
2022-02-15T09:58:09-0500

"t=\frac{t_0}{\sqrt{1-\frac{v^2}{c^2}}}=48^h."